NBA Youngboy correctly answered 85% of the questions on his last math test. If he missed 6 problems, how many questions were on the test?

Answer by ikleyn(53942)

.

    (1 - 0.85)x = 6

    0.15x       = 6

        x       = 6/0.15 = 40.     ANSWER

Solved.